OpenStudy (anonymous):

Solve the following inequality.6 − 4(x − 5) ≤ −2x + 12

OpenStudy (anonymous):

6-4x+20 <= -2x+12 4x+26-26 <= -2x+12-26 4x <= -2x-14 6x < = -14 x<=-7/3